OpenStudy (anonymous):

if f(a.b)=f(a)+f(b) and f(2)=3, then f(32) equals ?

hartnn (hartnn):

so, can you factor 32 in factors of 2 ?

OpenStudy (anonymous):

It could b either 4.8 or 2.16

hartnn (hartnn):

we have f(2) so we will select 2. 16 ok?

hartnn (hartnn):

so, f(32) = f(2*16) = f(2)+ f(16) got this ?

OpenStudy (anonymous):

but what about 16 ? Should we break it too ?

hartnn (hartnn):

absolutely!

hartnn (hartnn):

16= 2* what ?

OpenStudy (anonymous):

2*2*2*2 ?

hartnn (hartnn):

so, we go like this, step by step, f(32) = f(16) +f(2) but f(16) = f(8*2)= f(8)+f(2) so f(32)= f(8)+f(2)+ f(2) got this ? can u continue ?

OpenStudy (anonymous):

yeah. thanks alot :)

hartnn (hartnn):

welcome ^_^ tell me what u get as f(32) if you want me to verify

OpenStudy (anonymous):

15 ?

hartnn (hartnn):

yes, i get 15 too :) correct!

OpenStudy (anonymous):

\[ f(a\times b)=f(a)+f(b) \]Look familiar? \[ \ln(a\times b)=\ln(a)+\ln(b) \]

hartnn (hartnn):

and bdw, @MissKhan \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{P}}\color{purple}{\text{E}}\color{red}{\text{N}}\color{orange}{\text{S}}\color{#e6e600}{\text{T}}\color{green}{\text{U}}\color{blue}{\text{D}}\color{purple}{\text{Y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \]

OpenStudy (anonymous):

Yeah it looks familiar but am not allowed to use Calculator so ln is not appropriate.

OpenStudy (anonymous):

Actually its: \[ \log_2(32) = \log_2(2^5) = 5 \]

OpenStudy (anonymous):

@hartnn Thankyou. First experience to any site :)

OpenStudy (anonymous):

Basically \(f(x) = 3\log_2(x)\)

hartnn (hartnn):

nice! hope you enjoy learning and helping with openstudy! :) let me know if you need any help browsing this site....

OpenStudy (anonymous):

So how are we suppose to get 15 by applying log ? nd without using calculator ?

OpenStudy (anonymous):

@hartnn Suree:)

OpenStudy (anonymous):

Actually, I'm just trying to show you an interseting way of looking at it. That is all.

hartnn (hartnn):

wio just gave you an insight of what that function f could be, what we did is enough.

OpenStudy (anonymous):

Math is all about noticing pattersn.

OpenStudy (anonymous):

i dunno..

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

okay